Square root of -2x+3 (plus) the Square root of x+3 =3
Please HELP!!
Thanks,
Ashley LaFollette

Answer by stanbon(75887)

Sqrt(-2x+3)+ Sqrt(x+3) =3
Square both sides to get:
(-2x+3)+2[sqrt[(-2x+3)(x+3)]+(x+3)=9
-x+6-9 = 2sqrt[(-2x+3)(x+3)]
-(x+3)=2sqrt[(-2x+3)(x+3)]
Square both sides again to get:
x^2+6x+9=4[(-2x+3)(x+3)]
x^2+6x+9=4[-2x^2-3x+9]
x^2+6x+9=-8x^2-12x+36
9x^2+18x-27=0
x^2+2x-3=0
(x+3)(x-1)=0
x=-3 or x=1
Both answers check out.
Cheers,
Stan H.